OpenStudy (anonymous):

**********I FAN AND MEDAL!!!!********** A system of equations is shown below: x + y = 3 2x - y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

adding both equations 3x = 9 so x =3

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

@sourav_aich can you show me the steps?

OpenStudy (anonymous):

add both the equations

OpenStudy (anonymous):

x + y = 3 + 2x - y = 6 ___________ 2x + 7 = 6 I did this wrong didn't I

OpenStudy (anonymous):

@sourav_aich

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after adding you will get 3x = 9 so x =3

OpenStudy (anonymous):

Oh, how did you get that?

OpenStudy (anonymous):

just by adding

OpenStudy (anonymous):

x+2x = 3x y + (-y) =0 6 + 3 =9

OpenStudy (anonymous):

Oh thank you!